Using the supply function, you can get the quantity supplied if there is a

given price. Substitute the peso price in variable P and Multiply it by slope +10, the

available answer will be added to 0 (intercept). From here you will get the answer

10 quantity supplied. In the second example, the price of 5 is multiplied by 10. The

quantity supplied is 50.


1
Expert's answer
2021-11-18T14:49:11-0500

a)

The supply function

"Q_s = a+ bp"


Where

"Q_s -" Quantity supplied 

"a -" Factors influencing demand besides price (intercept)

"b" - Slope

"P" - Price


"Q_s =a+10*1"


but the intercept is


Therefore


"Q_s =0+10*1"


"Q_s =10"


b)

The supply function

"Q_s = a+ bp"


"Q_s =a+10*5"


but the intercept is


Therefore


"Q_s =0+10*5"


"Q_s =50"